A 2 m long and 3 m wide horizontal rectangular plate is submerged in water (at 20°C). The distance of the top surface of the plate from the free surface of the water is 5 m. The atmospheric pressure is 95 kPa. Considering atmospheric pressure, the hydrostatic force, in kN, acting on the top surface of this plate is

a. 307

b. 688

c. 747

d. 864

e. 2950
Answer #1

Atmospheric pressure= 95 KPa

Pressure due to water on plate= (unit weight of water*height)

=9.81*5

=49.05 KPa

Total pressure = 95+49.05

=144.05 Kpa

Force = pressure * Area of plate

= 144.05*6

=864.2 KN

(option D is correct)